an overhead electric power line carries a maximum current of 250 A. what is the magnitude of the maximum magnetic field in telsa at a point 0.037 m directly below the power line?
Answer #1

Solution)

Here,

The power line can be considered as an infinite wire.

So,

Magnetic field at a distance r away is calculated by:

B=uoI/2*3.14*r= (4*3.14*10^-7)*250 A/2*3.14*0.037

B= 0.00135 T

==========

===========